# The Unbounded Denominators Conjecture

### v3

@@ 6.2.5. A historical note
At this point we have established all the pieces for the proof of our main result. By Theorem 5.1.4, assumption (1) in Proposition 3.0.1 is indeed satisfied, with the sharp constant $A:=\zeta(3)/2>0$. By Theorem 6.0.1 with the choices $p(x):=x^{N}$ and $r:=1-BN^{-3}$, assumption (2) in Proposition 3.0.1 is also satisfied. In terms of the algebras of modular forms $M_{N}$ and $R_{N}$ at an even Wohlfahrt level $N$ introduced in 4.2.1, the conclusion of Proposition 3.0.1 is thus an inequality $R_{N}:M_{2}]\leq CN^{3}\log{N}$, for some absolute constant $C\in\mathbf{R}$ independent of $N$. At this point Proposition 4.3.5 proves the equality $R_{N}=M_{N}$ for all $N\in\mathbf{N}$, which is the unbounded denominators conjecture.
@@
Since $f(\tau)$ is a modular form, we are reduced to the situation of a number field $K$ such that $f(\tau)\in O_{K}\llbracket q^{1/N}\rrbracket$. We use $R_{2N}$ to denote the $K(\lambda)$-algebra generated by modular functions with coefficients in $K$, bounded denominators at $\zeta=i\infty$, and cusp widths dividing $2N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$ (similar to Definition 4.2.1). We follow the proof of Proposition 3.0.1 now on the case of the $K(\lambda)$-vector space $\mathcal{V}(U,x(t),O_{K})$ from Definition 2.0.4. Then $R_{2N}\subset\mathcal{V}(U,x(t),O_{K})$. Note that $U$ is stable under the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$, and thus $\mathcal{V}(U,x(t),O_{K})=\mathcal{V}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Q}}K$ and $\dim_{K(\lambda)}\mathcal{V}(U,x(t),O_{K})=\dim_{\mathbf{Q}(\lambda)}\mathcal{V}(U,x(t),\mathbf{Z})$. Thus by Corollary 2.0.5, we still have that $R_{2N}$ has dimension at most $CN^{3}\log N$ over $K(\lambda)$. The claimed extension to $\overline{\mathbf{Z}}\llbracket q^{1/N}\rrbracket$ Fourier expansions now follows upon remarking that the proof of Proposition 4.3.5 still persists when $\mathbf{Q}$ is replaced by $K$.
@@ 7.1. Generalized McKay–Thompson series with roots from Monstrous Moonshine
Our argument also proves a vector generalization of the unbounded denominators conjecture, which has been conjectured by Mason [Mas12] (see also the earlier work of Kohnen and Mason [KM08] for a special case) to the setting of vector-valued modular forms of $\mathrm{SL}_{2}(\mathbf{Z})$, with motivation from the theory of vertex operator algebras and the Monstrous Moonshine conjectures. The weaker statement of algebraicity over the ring of modular forms was conjectured earlier by Anderson and Moore [AM88], within the context of the partition functions or McKay–Thompson series attached to rational conformal field theories. We refer also to André [And04, Appendix], for a discussion from the arithmetic algebraization point of view — the method that we build upon in our present paper — on the Grothendieck–Katz $p$-curvature conjecture. Eventually the more precise expectation crystallized, see Eholzer [Eho95, Conjecture on page 628], that all RCFT graded twisted characters are in fact classical modular forms for a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ (which is more precise than Anderson and Moore’s conjectured algebraicity over the modular ring $\mathbf{Z}[E_{4},E_{6}]$).
@@
Our paper, via Theorem 7.3.3 below for the vector valued extension of the congruence property, inherits a new proof of these modularity theorems. The connection was engineered by Knopp and Mason [KM03a], with their formalization of generalized modular forms for $\mathrm{SL}_{2}(\mathbf{Z})$, and fine tuned by Kohnen and Mason [KM08, § 4], who brought forward the idea of a purely arithmetic approach — based on the integrality properties of the Fourier coefficients, that record a graded dimension and are hence integers — for a part of Borcherds’s theorem [Bor92] (the Conway–Norton “Monstrous Moonshine” conjecture). Namely, suppressing the Hauptmodul property, for the classical modularity — under a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ — of all the various McKay–Thompson series for the Monster group over the Moonshine module $V^{\sharp}$. Whereas Borcherds’s proof, based on his own generalized Kac–Moody algebras that go outside of the general framework of vertex operator algebras, is rather particular to the Monster vertex algebra and genus 0 arithmetic groups, Kohnen and Mason proposed that an arithmetic abstraction from the integrality of Fourier coefficients might open up a window on the modularity and congruence properties to apply just as well in the equivariant setting to any rational $C_{2}$-cofinite vertex operator algebra — this theorem, eventually proved in [DLN15, DR18] by other means, was an open problem at the time of [KM08].
@@ 7.2. Unbounded denominators for the solutions of certain ODEs
Our condition is that the $x=0$ local monodromy group is $\mathbf{Z}/N$ for some $N\in\mathbf{N}$. Then the formal function $g(x):=f(x^{N})$ is in $\mathbf{Z}\llbracket x\rrbracket$ and fulfills a linear ODE on $\mathbf{P}^{1}\smallsetminus\{16^{-1/N}\mu_{N},\infty\}$. In our notation of Theorem 2.0.5, that means $g\in\mathcal{H}(\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N},\mathbf{Z})$. Hence, denoting again by $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ the universal covering map taking $F_{N}(0)=0$, recalling our exact uniformization radius formula in Theorem 5.1.4 giving in particular the strict lower bound
@@
for some parameter $r$ with $\sqrt[N]{16}\big/|F_{N}^{\prime}(0)|<r<1$, Theorem 2.0.5 implies that $g(x)\in\mathbf{Z}\llbracket x\rrbracket$ is an algebraic power series. Hence $f(x)=g(\sqrt[N]{x})$ is algebraic.
@@
It remains to us an open question whether a complete description of all integral solutions $f\in\mathbf{Z}\llbracket x\rrbracket$ on dropping the $x=0$ finite local monodromy condition in Theorem 7.2.1 should arise in this way from a classical congruence modular form expressed into a holonomic function in $x=\lambda/16$. We formulate the precise statement in Question 7.4.3 below.
@@ 7.3. Vector-valued modular forms
We close our paper by another formulation of Theorem 7.2.1, translated now over to the language of vector-valued modular forms. The following definition is a special case of vector-valued modular forms studied in [FM16a, §2, Definition 1]
@@
A vector-valued modular form of weight $k\in\mathbf{Z}$ and dimension $n$ for $\mathrm{SL}_{2}(\mathbf{Z})$ is a pair $(F,\rho)$ made of a holomorphic mapping $F=(F_{1},\ldots,F_{n}):\mathbf{H}\to\mathbf{C}^{n}$ and an $n$-dimensional complex representation
$\rho:\mathrm{SL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$
@@
  • For all $\gamma\in\mathrm{SL}_{2}(\mathbf{Z})$,
    $F^{\mathrm{t}}\,|_{k}\gamma=\rho(\gamma)F^{\mathrm{t}};$
  • The matrix
    $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\in\mathrm{GL}_{n}(\mathbf{C})$
    is semisimple.
@@
    $\forall\tau\in\mathbf{H},\quad a\leq\mathrm{Re}\,\tau\leq b,\quad\mathrm{Im}\,\tau\geq C\quad\Longrightarrow\quad|F_{j}(\tau)|\leq Ae^{B\,\mathrm{Im}\,\tau}.$
@@
Taken together, see [AM88, § 2.A], the semisimplicity and moderate growth conditions are equivalent to the existence of generalized Puiseux formal expansions (except in general with irrational exponents: but without $\log{q}$ terms, due to semisimplicity) of each component function $F_{j}(\tau)$ at the cusp $q=0$. More precisely, via a change of basis (see the equivalent notion in [FM16a]), we may assume that $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ is a diagonal matrix. If $F_{j}$ is a $\lambda$-eigenvector of $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$, then $F_{j}=\sum_{n\in\mathbf{Z}_{\geq n_{0}}}a_{n,j}q^{n+\mu}$ for some $n_{0}\in\mathbf{Z}$, where $q=e^{2\pi i\tau}$ and we choose a $\mu\in\mathbf{C}$ such that $\lambda=e^{2\pi i\mu}$.
@@
Thus, the classical (scalar-valued) modular forms $M_{k}(\Gamma(1),\chi)$ attached to a finite-order character $\chi:\Gamma(1)\to U(1)$ are precisely the special case $n=1$ of one-dimensional vector-valued modular forms and a unitary character $\rho$. In a reverse direction, any classical (scalar-valued) modular form for a finite index subgroup $\Gamma\subseteq\mathrm{SL}_{2}(\mathbf{Z})$ can be considered as the first component of a vector-valued modular form for $\mathrm{SL}_{2}(\mathbf{Z})$ of dimension $\Gamma(1):\Gamma]$. From that point of view, there is no loss of generality in Definition 7.3.1 to limit to the representations of the ambient group $\mathrm{SL}_{2}(\mathbf{Z})$.
@@
Knopp and Mason’s generalized modular forms [KM03a] are the case, intermediate in generality, where the representation $\rho$ is monomial: that is, induced from a linear character $\chi:\Gamma\to\mathbf{C}^{\times}$ on a finite index subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$. If that character $\chi$ is unitary, then in fact it has finite image and all components of $F$ are classical modular forms of weight $k$ for a finite index subgroup [KM03a]. The general (non-unitary) case does come up for the partition function and correlation functions of a rational conformal field theory [KM03a], to which the point of contact is supplied by Zhu’s modularity theorem [Zhu96] (see also Codogni [Cod20] for a recent different proof and a generalization), and its extension to the equivariant setting by Dong, Li and Mason [DLM00].
To make the connection to Theorem 7.2.1, note upon restricting the representation $\rho$ to the free subgroup $\mathbf{Z}\ast\mathbf{Z}\cong\Gamma(2)\subset\Gamma(1)=\mathrm{SL}_{2}(\mathbf{Z})$ that the case of weight $k=0$ and finite-order element $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ is equivalent to exactly the situation of 7.2.1: a local system on the triply-punctured projective line $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ that has a finite local monodromy around the puncture $x=0$. We refer the reader to [BG07, Gan14] regarding the bridge between these two equivalent points of view.
@@
Let $(F,\rho)$ be a vector-valued modular form for $\mathrm{SL}_{2}(\mathbf{Z})$ of dimension $n$ and weight $k$. Suppose that some component function $F_{j}(\tau):\mathbf{H}\to\mathbf{C}$ of $F=(F_{1},\ldots,F_{n}):\mathbf{H}\to\mathbf{C}^{n}$ has at $\tau=i\infty$ a formal Fourier expansion lying in $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket e^{\pi i\tau}\rrbracket$. Then that component $F_{j}(\tau)$ is a classical modular form of weight $k$ on a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$.
@@
After some standard theorems from the theory of $G$-functions to reduce to the case that the semisimple matrix $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\in\mathrm{GL}_{n}(\mathbf{C})$ is in fact of finite order, this is an equivalent expression of Theorem 7.2.1.
The transition is as follows. By taking the componentwise product $F(\tau)g(\tau)(\lambda(\tau)/16\Delta(\tau/2))^{\frac{k+k^{\prime}}{12}}$, where we choose a non-zero scalar-valued modular form $g(\tau)\in\mathbf{Z}[[q]]$ of weight $k^{\prime}$ such that $12\mid k+k^{\prime}$, and restricting the full modular group to its subgroup $\Gamma(2)$, we reduce to the case $k=0$ of local systems on $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$. Without loss of generality upon passing to a factor, we may assume that local system to be irreducible. The holomorphic vector bundle with integrable connection admitting $F$ for its horizontal sections is indeed meromorphic at the cusps of $Y(2)$ due to the existence of the $q$-expansion of $F$. Hence $F$ is a solution of a rank-$n$ system of first-order linear homogeneous ODEs over $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$. By the theorem of the cyclic vector, see [DGS94, § III.4], there is an irreducible linear differential operator $L$ over $\mathbf{Q}(\lambda)$ without singularities on $Y(2)=\mathrm{Spec}\,\mathbf{Z}[\lambda,1/\lambda,1/(1-\lambda)]$ and such that all $n$ component functions $F_{1},\ldots,F_{n}$ are formal solutions of the linear homogeneous ODE $L(f)=0$. Since one of these (namely, $F_{j}$) has a $\lambda=0$ formal expansion in $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket\lambda/16\rrbracket$, Chudnovsky’s theorem [DGS94, Theorem VIII.1.5] implies that $L$ satisfies the Galočkin (finite global operator height $\sigma(L)<\infty$) condition [DGS94, VII.2.(2.3) on page 227], hence by the Bombieri–André theorem [DGS94, Theorem VII.2.1], $L$ satisfies the Bombieri (finite generic global inverse radius $\rho(L)<\infty$) condition [DGS94, VII.2.(2.1) on page 226], and is therefore globally nilpotent. At this point Katz’s local monodromy theorem [Kat70] (see also [DGS94, Theorem III 2.3 (ii)]) proves that $L$ has quasi-unipotent local monodromies. Now by the semisimplicity condition on $\rho\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$ in Definition 7.3.1, it follows that in fact the $x=0$ local monodromy of $L$ has finite order. The result then follows on applying Theorem 7.2.1 to $f(x)=F_{j}(\tau)$. ∎
@@
If all components of a vector-valued modular form $(F,\rho)$ for $\mathrm{SL}_{2}(\mathbf{Z})$ have Fourier expansions with bounded denominators, then the representation $\rho$ has a finite image, and more precisely $\ker(\rho)\supseteq\Gamma(N)$ for some $N\in\mathbf{N}$.

### v4

@@ 6.2.17. A historical note
At this point we have established all the pieces for the proof of our main result. By Theorem 5.1.4, assumption (1) in Proposition 3.0.1 is indeed satisfied, with the sharp constant $A:=\zeta(3)/2>0$. By Theorem 6.0.1 with the choices $p(x):=x^{N}$ and $r:=1-BN^{-3}$, assumption (2) in Proposition 3.0.1 is also satisfied. In terms of the algebras of modular forms $M_{2N}$ and $R_{2N}$ at an even Wohlfahrt level $2N$ introduced in 4.2.1, the conclusion of Proposition 3.0.1 is thus an inequality $R_{2N}:M_{2}]\leq CN^{3}\log{N}$, for some absolute implicit constant $C\in\mathbf{R}$ independent of $N$. At this point Proposition 4.3.5 proves the equality $R_{2N}=M_{2N}$ for all $N\in\mathbf{N}_{>0}$, which is the unbounded denominators conjecture.
@@
Since $f(\tau)$ is a modular form, we are reduced to the situation of a number field $K$ such that $f(\tau)\in O_{K}\llbracket q^{1/N}\rrbracket$. We use $R_{2N}$ to denote the $K(\lambda)$-algebra generated by modular functions with coefficients in $K$, bounded denominators at $\zeta=i\infty$, and cusp widths dividing $2N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$ (similar to Definition 4.2.1). We follow the proof of Proposition 3.0.1 now on the case of the $K(\lambda)$-vector space $\mathcal{V}(U,x(t),O_{K})$ from Definition 2.0.4. Then $R_{2N}\subset\mathcal{V}(U,x(t),O_{K})$. Note that $U$ is stable under the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$, and thus $\mathcal{V}(U,x(t),O_{K})=\mathcal{V}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Q}}K$ and $\dim_{K(\lambda)}\mathcal{V}(U,x(t),O_{K})=\dim_{\mathbf{Q}(\lambda)}\mathcal{V}(U,x(t),\mathbf{Z})$. Thus by Corollary 2.0.5, Theorem 5.1.4, and Theorem 6.0.1, we still have that $R_{2N}$ has dimension at most $CN^{3}\log N$ over $K(\lambda)$. The claimed extension to $\overline{\mathbf{Z}}\llbracket q^{1/N}\rrbracket$ Fourier expansions now follows upon remarking that the proof of Proposition 4.3.5 still persists when $\mathbf{Q}$ is replaced by $K$.
@@ 7.1. Generalized McKay–Thompson series with roots from monstrous moonshine
Our argument also proves a vector generalization of the unbounded denominators conjecture, which has been conjectured by Mason [Mas12] (see also the earlier work of Kohnen and Mason [KM08] for a special case) to the setting of vector-valued modular forms of $\mathrm{SL}_{2}(\mathbf{Z})$, with motivation from the theory of vertex operator algebras and the monstrous moonshine conjectures. The weaker statement of algebraicity over the ring of modular forms was conjectured earlier by Anderson and Moore [AM88], within the context of the partition functions or McKay–Thompson series attached to rational conformal field theories. We refer also to André [And04, Appendix] for a discussion from the arithmetic algebraization point of view — the method that we build upon in our present paper — on the Grothendieck–Katz $p$-curvature conjecture. Eventually the more precise expectation crystallized (see Eholzer [Eho95, Conjecture on page 628]) that all rational conformal field theory graded twisted characters are in fact classical modular forms for a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$, which is more precise than Anderson and Moore’s conjectured algebraicity over the modular ring $\mathbf{Z}[E_{4},E_{6}]$.
@@
Our paper, via Theorem 7.3.3 below for the vector valued extension of the congruence property, inherits a new proof of these modularity theorems. The connection was engineered by Knopp and Mason [KM03a], with their formalization of generalized modular forms for $\mathrm{SL}_{2}(\mathbf{Z})$, and fine tuned by Kohnen and Mason [KM08, § 4], who brought forward the idea of a purely arithmetic approach — based on the integrality properties of the Fourier coefficients, that record a graded dimension and are hence integers — for a part of Borcherds’ theorem [Bor92] (the Conway–Norton “monstrous moonshine” conjecture). Namely, suppressing the Hauptmodul property, for the classical modularity — under a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ — of all the various McKay–Thompson series for the Monster group over the moonshine module $V^{\sharp}$. Whereas Borcherds’ proof, based on his own generalized Kac–Moody algebras that go outside of the general framework of vertex operator algebras, is rather particular to the Monster vertex algebra and genus 0 arithmetic groups, Kohnen and Mason proposed that an arithmetic abstraction from the integrality of Fourier coefficients might open up a window on the modularity and congruence properties to apply just as well in the equivariant setting to any rational $C_{2}$-cofinite vertex operator algebra — this theorem, eventually proved in [DLN15, DR18] by other means, was an open problem at the time of [KM08].
@@ 7.2. Unbounded denominators for the solutions of certain ODEs
Our condition is that the $x=0$ local monodromy group is $\mathbf{Z}/N$ for some $N\in\mathbf{N}_{>0}$. Then the formal function $g(x):=f(x^{N})$ is in $\mathbf{Z}\llbracket x\rrbracket$ and fulfills a linear ODE on $\mathbf{P}^{1}\smallsetminus\{16^{-1/N}\mu_{N},\infty\}$. In our notation of Corollary 2.0.5, that means $g\in\mathcal{H}(\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N},\mathbf{Z})$. Hence, denoting again by $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ the universal covering map taking $F_{N}(0)=0$, recalling our exact uniformization radius formula in Theorem 5.1.4 giving in particular the strict lower bound
@@
for some parameter $r$ with $\sqrt[N]{16}\big/|F_{N}^{\prime}(0)|<r<1$, Corollary 2.0.5 implies that $g(x)\in\mathbf{Z}\llbracket x\rrbracket$ is an algebraic power series. Hence $f(x)=g(\sqrt[N]{x})$ is algebraic.
@@
It remains to us an open question whether a complete description of all integral solutions $f\in\mathbf{Z}\llbracket x\rrbracket$ on dropping the $x=0$ finite local monodromy condition in Theorem 7.2.1 should arise in this way from a classical congruence modular form expressed into a holonomic function in $x=\lambda/16$. We formulate the precise statement in Question 7.4.5 below.
@@ 7.3. Vector-valued modular forms
We close our paper by another formulation of Theorem 7.2.1, translated now over to the language of vector-valued modular forms. The following definition is a special case of the vector-valued modular forms studied in [FM16a, § 2].
@@
A vector-valued modular form of weight $k\in\mathbf{Z}$ and dimension $n$ for $\mathrm{PSL}_{2}(\mathbf{Z})$ is a pair $(F,\rho)$ made of a holomorphic mapping $F=(F_{1},\ldots,F_{n}):\mathbf{H}\to\mathbf{C}^{n}$ and an $n$-dimensional complex representation
$\rho:\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$
@@
  • For all $\gamma\in\mathrm{PSL}_{2}(\mathbf{Z})$, we have $F^{\mathrm{t}}\,|_{k}\gamma=\rho(\gamma)F^{\mathrm{t}}$.
  • The matrix $\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)\in\mathrm{GL}_{n}(\mathbf{C})$ is semisimple.
@@
    $\tau\in\mathbf{H},\quad a\leq\mathrm{Re}\,\tau\leq b,\quad\mathrm{Im}\,\tau\geq C\quad\Longrightarrow\quad|F_{j}(\tau)|\leq Ae^{B\,\mathrm{Im}\,\tau}.$
@@
Taken together (see, for example, [AM88, § 2.A]) the semisimplicity and moderate growth conditions are equivalent to the existence of generalized Puiseux formal expansions (except in general with irrational exponents: but without $\log{q}$ terms, due to semisimplicity) of each component function $F_{j}(\tau)$ at the cusp $q=0$. More precisely, via a change of basis (see the equivalent notion in [FM16a]), we may assume that $\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)$ is a diagonal matrix. If $F_{j}$ is a $\lambda$-eigenvector of $\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)$, then $F_{j}=\sum_{n\in\mathbf{Z}_{\geq n_{0}}}a_{n,j}q^{n+\mu}$ for some $n_{0}\in\mathbf{Z}$, where $q=e^{2\pi i\tau}$ and we choose a $\mu\in\mathbf{C}$ such that $\lambda=e^{2\pi i\mu}$.
@@
Thus, the classical (scalar-valued) modular forms $M_{k}(\Gamma(1),\chi)$ attached to a finite-order character $\chi:\Gamma(1)\to U(1)$ are precisely the special case $n=1$ of one-dimensional vector-valued modular forms and a unitary character $\rho$. In a reverse direction, any classical (scalar-valued) modular form for a finite index subgroup $\Gamma\subseteq\mathrm{PSL}_{2}(\mathbf{Z})$ can be considered as the first component of a vector-valued modular form for $\mathrm{PSL}_{2}(\mathbf{Z})$ of dimension $\Gamma(1):\Gamma]$. From that point of view, there is no loss of generality in Definition 7.3.1 to limit to the representations of the ambient group $\mathrm{PSL}_{2}(\mathbf{Z})$. Here and in the following, we continue to denote by $\rho$ the extended homomorphism $\rho:\Gamma(1)=\mathrm{SL}_{2}(\mathbf{Z})\to\mathrm{PSL}_{2}(\mathbf{Z})\to\mathrm{GL}_{n}(\mathbf{C})$, which (by convention) contains $Z(\mathrm{SL}_{2}(\mathbf{Z}))=\{\pm I\}$ in its kernel.
@@
Knopp and Mason’s generalized modular forms [KM03a] are the case, intermediate in generality, where the representation $\rho$ is monomial: that is, induced from a linear character $\chi:\Gamma\to\mathbf{C}^{\times}$ on a finite index subgroup $\Gamma\subset\mathrm{PSL}_{2}(\mathbf{Z})$. If that character $\chi$ is unitary, then in fact it has finite image and all components of $F$ are classical modular forms of weight $k$ for a finite index subgroup [KM03a]. The general (non-unitary) case does come up for the partition function and correlation functions of a rational conformal field theory [KM03a], to which the point of contact is supplied by Zhu’s modularity theorem [Zhu96], and its extension to the equivariant setting by Dong, Li and Mason [DLM00].
To make the connection to Theorem 7.2.1, note upon restricting the representation $\rho$ to the free subgroup
$\mathbf{Z}\ast\mathbf{Z}=\left<\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\begin{pmatrix}1&0\\ 2&1\end{pmatrix}\right>\subset\Gamma(2)\subset\Gamma(1)=\mathrm{SL}_{2}(\mathbf{Z})$
that the case of weight $k=0$ and finite-order element $\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)$ is equivalent to exactly the situation of 7.2.1: a local system on the triply-punctured projective line $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ that has a finite local monodromy around the puncture $x=0$. (Note that the map from $\mathbf{Z}\ast\mathbf{Z}$ to $\Gamma(2)/\pm I\subset\mathrm{PSL}_{2}(\mathbf{Z})$ is an isomorphism.) Concretely, the local system with integrable connection $(\mathcal{E},\nabla)$ over $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is defined by taking for $\nabla$ the derivation $d/dx$ in the coordinate $x:=\lambda/16$ of the base curve $Y(2)$, and the vector bundle $\mathcal{E}\to Y(2)$ over the base algebraic curve $Y(2)=\mathrm{Spec}\,\mathbf{Z}[x,16/x,1/(1-16x)]$ to be defined by the rank-$6n$ free $\mathbf{Z}[x,1/x,1/(1-16x)]$-module spanned by the functions $F_{j}|\gamma$, where $F_{1},\ldots,F_{n}$ range over the components of the vector-valued modular form $F$ on the modular group $\Gamma(1)=\mathrm{SL}_{2}(\mathbf{Z})$, and $\gamma$ runs through all six cosets for $\Gamma(2)$ in $\Gamma(1)$ (with the stroke action here being in weight 0). The multiplier system $\rho$ features as the monodromy representation $\rho|_{\Gamma(2)}:\Gamma(2)\to\mathrm{GL}_{n}(\mathbf{C})$. The condition in Remark 7.3.2 on the existence of a Fourier expansion at the cusp of $\mathrm{PSL}_{2}(\mathbf{Z})$ means that the local system with integrable connection $(\mathcal{E},\nabla)$ has regular singularities with semisimple local monodromies around the three cusps of $Y(2)$. We refer the reader to [BG07] and [Gan14] for the bridge between these two equivalent points of view.
@@
Let $(F,\rho)$ be a vector-valued modular form for $\mathrm{PSL}_{2}(\mathbf{Z})$ of dimension $n$ and weight $k$. Suppose that some component function $F_{j}(\tau):\mathbf{H}\to\mathbf{C}$ of $F=(F_{1},\ldots,F_{n}):\mathbf{H}\to\mathbf{C}^{n}$ has at $\tau=i\infty$ a formal Fourier expansion lying in $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket e^{\pi i\tau}\rrbracket$. Then that component $F_{j}(\tau)$ is a classical modular form of weight $k$ on a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$.
@@
After some standard theorems from the theory of $G$-functions to reduce to the case that the semisimple matrix $\rho\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)\in\mathrm{GL}_{n}(\mathbf{C})$ is in fact of finite order, this is an equivalent expression of Theorem 7.2.1. The transition is as follows. First, by taking the componentwise product
$F(\tau)g(\tau)\left(\frac{\lambda(\tau)}{16\Delta(\tau/2)}\right)^{\frac{k+k^{\prime}}{12}},$
where we choose a non-zero scalar-valued modular form $g(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ of weight $k^{\prime}$ to have $12\mid k+k^{\prime}$, we reduce to the case of a vector-valued modular form of weight $k=0$ on $\mathrm{PSL}_{2}(\mathbf{Z})$.
We restrict that form from $\mathrm{PSL}_{2}(\mathbf{Z})$ to its index-$6$ torsion-free subgroup $\Gamma(2)/\langle\pm I\rangle$. The remarks immediately preceding the statement of the theorem-under-proof construct a rank $6n$ local system with integrable connection $(\mathcal{E},\nabla)$ over the modular curve $Y(2)=\mathrm{Spec}\,{\mathbf{Z}[x,16/x,1/(1-16x)}]$. The monodromy representation of that local system is the homomorphism $\rho|_{\Gamma(2)}:\Gamma(2)\to\mathrm{GL}_{n}(\mathbf{C})$. As we study the specific component function $F_{j}$ and its monodromy unfoldings in this local system, upon replacing the range $\mathrm{GL}_{n}(\mathbf{C})$ by a lower-dimensional general linear group, we lose no generality in assuming that the local system is irreducible.
With these reductions, we have realized our original $\mathrm{PSL}_{2}(\mathbf{Z})$ vector-valued modular form component $F_{j}\in\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket\lambda/16\rrbracket$ of interest as one of the $6n$ component power series in a complex Puiseux series vector solution to an irreducible rank-$6n$ system of first-order linear homogeneous ODEs over $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$. One of the components — namely, $F_{j}$ — in the solution vector to this irreducible linear differential system is a $G$-function [DGS94, page xiii]. David and Gregory Chudnovsky’s fundamental $G$-functions theorem [DGS94, Theorem VIII.1.5] implies that this linear differential system $(\mathcal{E},\nabla)$ satisfies the Galočkin (finite global operator height $\sigma(\nabla)<\infty$) condition [DGS94, VII.2.(2.3) on page 227], hence by the Bombieri–André theorem [DGS94, Theorem VII.2.1], it satisfies the Bombieri (finite generic global inverse radius $\rho(\nabla)<\infty$) condition [DGS94, VII.2.(2.1) on page 226], and is therefore globally nilpotent. At this point Katz’s local monodromy theorem [Kat70] (see also [DGS94, Theorem III 2.3 (ii)]) proves that $(\mathcal{E},\nabla)$ has quasi-unipotent local monodromies. Since, as we already observed from Remark 7.3.2, our local system $(\mathcal{E},\nabla)$ also has semisimple local monodromies, it follows that these local monodromies have finite order.
Thus we find that $f:=F_{j}\in\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ (in the coordinate $x:=\lambda/16$) satisfies a linear ODE $L(f)=0$ over $Y(2)=\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ with a finite local monodromy at $x=0$. The result now follows on applying Theorem 7.2.1 to $f(x)=F_{j}(\tau)$. ∎
@@
If all components of a vector-valued modular form $(F,\rho)$ for $\mathrm{PSL}_{2}(\mathbf{Z})$ have Fourier expansions with bounded denominators, then the representation $\rho$ has a finite image, and more precisely $\ker(\rho)\supseteq\Gamma(N)$ for some $N\in\mathbf{N}_{>0}$.